Show that {fn}, where fn : [0, 1] → R is the map fn(x) = x
n, is not uniformly convergent.
1
Expert's answer
2020-06-22T18:07:34-0400

Given that , "(f_n)" be a sequence of function.

Where "f_n:[0,1]\rightarrow\R" defined by "f_n(x)=x^n" .

Clearly, if "x=1" then the sequence "(f_n(1))" converges to 1 .

If "0\leq x <1" then the sequence "(f_n(x))" converges to 0 as we known that "lim_{n\to \infty} x^n=0" if "0\leq x<1" .

Let "f(x)=\begin{cases}
1 & \text{if} \ x=1 \\
0 & \text{if} \ 0\leq x <1
\end{cases}"

Thus the sequence of function "(f_n(x))" convergent to "f" on the set "[0,1]" .

If "n_k=k \ and \ x_k=(\frac{1}{2})^{ \frac{1}{k}}" then

"| f_{n_k}(x_k)-f(x_k)|=|\frac{1}{2}-0|=\frac{1}{2}" .


Therefore the sequence "(f_k)" doesn't converge uniformly on "[0,1] \ to \ f" .